OpenStudy (anonymous):

A stone is dropped from the top of a tower of height 40m. The stone falls from rest and air resistance is negligible. What is the time taken for it to fall the last 10m to the ground?

OpenStudy (anonymous):

Ok I assumed: a=9.81 u=0 s=10m And using the equation s= ut + 1/2 at2 I calculated the time to be 2.85--2.9 seconds. Buuuuut my answer is wrong!! Did i assume something wrong?

OpenStudy (anonymous):

Use S=Ut+1/2at^2 40=0t+1/2*9.81t^2 40=4.905t^2 8.15=t^2 t=2.86 that is the total time for the stone to fall 40m now if we divide the time by 4 since 40/4=10m we get a time of .71 seconds

OpenStudy (anonymous):

does this make sense at all?

OpenStudy (anonymous):

It does, but the answer is 0.38s !! Idk maybe i'm mindblocked, idk how to get thisssssssss. :/

OpenStudy (anonymous):

ok i think i got it! if u figure out how long it take to fall 30m and figure out how long it takes for the stone to fall 40m all you have to do is subtract the 40m time- 30m time and you will get .38s

OpenStudy (anonymous):

2.86-2.47=.38s

OpenStudy (anonymous):

Vi=0 H=40m h=H-10m=30m g=9.8m/*s^2 As we know that 2gh=Vf^2-Vi^2 2*9.8*30=Vf^2 Vf=24.2487m/s now as ds/dt=Vf=24.2487 ds=24.2487dt integratin between \[0\le s \le10m \] we get 10m=24.2487t t=10/24.2487=0.40 sec

OpenStudy (anonymous):

@Brent0423 ohhhhhhhhhhhhhhhhhhhhhhhhhh!!! Gotcha. Thanks a bunch!